Sibling IQ Scores: There have been numerous studies involving the correlation and differences in IQ's among siblings. Here we consider a small example of such a study. We will test the claim that, on average, older siblings have a higher IQ than their younger sibling. The results are depicted for a sample of 10siblings in the table below. Test the claim at the 0.01 significance level. You may assume the sample of differences comes from a normally distributed population.

Pair ID Older Sibling IQ (x) Younger Sibling IQ(y) difference (d = xy)   
1 86 84 2
2 89 92 -3
3 93 89 4
4 94 94 0
5 101 97 4
6 106 104 2
7 107 107 0
8 112 111 1
9 116 110 6
10 123 115 8
Mean 102.7 100.3 2.4
s 12.2 10.5 3.2

If you are using software, you should be able copy and paste the data directly into your software program.

(a) The claim is that the mean difference is positive (μd > 0). What type of test is this?

This is a right-tailed test.This is a two-tailed test.     This is a left-tailed test.


(b) What is the test statistic? Round your answer to 2 decimal places.
td =  
To account for hand calculations -vs- software, your answer must be within 0.01 of the true answer.

(c) Use software to get the P-value of the test statistic. Round to 4 decimal places.
P-value =
Answer #1

Data Summary

n Mean (M) Variance (SS) Standard Deviation (SD)
Older Sibling IQ 10 102.7 149.3445 12.2207
Younger Sibling IQ 10 100.3 110.6778 10.5204

a)   The null and alternative hypotheses are                      
   Ho : μd ≤ 0       μd is the difference μ1 - μ2              
   Ha : μd > 0                      
   This is a right-tailed test                     
                          
                          
b)   We use independent Samples T test since population standard deviation is unknown, and sample size is small                      
   Using the formulae given below, we get                      
                          
   Degrees of Freedom                      

df1 = n1 - 1   df2 = n2 - 1   df = n1 + n2 - 2

    df1 = 9   df2 = 9                  
   df = 18                      
                          
   Pooled Variance Sp²  

                   
   Sp² = 130.0111                      
                           
   Mean Squared Error S(M1-M2)    

               
   S(M1-M2) = 5.0992                      
                          
   t-statistic  

                  
   td = 0.47                      
                          
                          
c)   For t = 0.47, df = 18 we find the Right Tailed p-value using Excel function t.dist                      
   p-value = 1 - t.dist(0.4707, 18, TRUE)                      
   p-value = 0.3218                      
                          
   Let the significance level α = 0.05                      
   Decision                      
   0.3218 > 0.05                      
   that is p-value > α                      
   Hence we DO NOT REJECT Ho                      
                          
   Conclusion                      
   There does not exist enough statistical evidence at α = 0.05 to show that                       
   the mean difference is positive                       
   In other words,                      
   older siblings need not have higher IQ than their younger siblings